find the first term of the arithmetic progression whose third term a3 is 7 and
whose eighth term a8 is 17
a)0 b) 2 c)3 d)1 e) 4


Answer by stanbon(75887)

find the first term of the arithmetic progression whose third term a3 is 7 and
whose eighth term a8 is 17
a)0 b) 2 c)3 d)1 e) 4
---------------------------------------
a(3) = a(1) + 2d = 7
a(8) = a(1) + 7d = 17
--------------------------------
Subtract a(3) from a(8) to get:
5d = 10
d = 2
---------------------
Substitute that into a(1) + 2d = 7 to solve for a(1)
a(1) + 2*2 = 7
a(1) = 3
==================
Cheers,
Stan H.
Reply to stanbon@comcast.net